OpenStudy (anonymous):

are these correct answers to these two prolbems any1? 14-2y-8y<94 i came out with the answer >-5 now for the second one its find the slope x+3y=6 i came out with the answer (2,6) are those right any1

OpenStudy (aroub):

for the first i guess ur answer is incorrect! :( i believe its y>-8 for the second one the slope is -1/3

OpenStudy (anonymous):

1st one 14-10y<94 10y>-80 y>-8 slope 4 2nd one is y=-x/3+2 m=-1/3

OpenStudy (anonymous):

thank you im so not good at math

OpenStudy (aroub):

its okay we are here to help...!! :)